Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 166     Entry time: 14 Sep 2024, 01:33
Author: Peter Zalden 
Group: FXE 
Subject: Status 

FXE received the beam and did not change accelerator settings.

Beam alignment to FXE was re-established using the new values for the old M1/M2 offset.

Monochromator aligned to the SPB energy setpoint.

THz source aligned to Xrays and several reference positions for Xray pointing taken. Further steps require THz radiation, which is not available at this point.

Xrays realigned back to SPB XTD9 screen.

ELOG V3.1.4-7c3fd00